# A Survey on Replication Studies and More Directions for Reasoning Language Models

### v1

@@ 1 Introduction
Reasoning language models (RLMs) have emerged as a transformative advancement in the evolution of large language models (LLMs), such as OpenAI o-series (Jaech et al., 2024; OpenAI, 2025), DeepSeek-R1 (Guo et al., 2025), and QwQ series (Team, 2024; 2025c). Unlike conventional LLMs that merely generate unstructured responses, these models incorporate an explicit chain-of-thought process, providing step-by-step reasoning that mimics human cognitive processes–such as invoking self-verification, reflection, and more. This shift quickly attracted attention of the LLM research community, as it meets the growing demand for better explainability in complex tasks like mathematical problem solving, code generation, and logical reasoning, as well as the pursuit of steadily increasing accuracy. The significance of RLMs lies in their potential to enhance the accuracy of language models’ response with trustful rationales. More than only providing answers, these models also reveal their thought process, which is invaluable for educational purposes, critical decision-making, and debugging AI reasoning errors. DeepSeek-R1 (Guo et al., 2025) is a prime example of this new generation. It leverages innovative training techniques such as supervised fine-tuning (SFT) and reinforcement learning from verifiable rewards (RLVR), allowing the model to develop powerful reasoning behaviors with affordable amount of supervised data. Especially, by using methods like RLVR and incorporating cold-start data, it achieves performance comparable to prior models with relatively low training costs.
However, many critical details remain undisclosed which is required to replicate the reasoning performance and the behaviors of self-verification and reflection exhibited by DeepSeek-R1. Although DeepSeek-R1 (Guo et al., 2025) have publicly released their solution as training the model with Group Relative Policy Optimization (GRPO, Shao et al. (2024)) and rule-based reward systems, the optimal design of the reinforcement learning algorithm and reward system remains underexplored. Moreover, the training data and configurations of the SFT and RLVR stages are not released, leaving the impact on model performance to be further examined. In response, many replication works have attempted to explore the optimal design for RLMs from various perspectives (see Figure 1), yet a comprehensive list and comparison of these works are still lacking.
@@
forked edges, for tree= grow=east, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, align=center, fill=theme!20, , [Training Reasoning Models, ver [Supervised Fine-tuning (§2), level1, fill=SFT1 [Supervised Fine-tuning Datasets (§2.1), level2,fill=SFT2 [e.g., OpenThoughts-114k(Team, 2025a), OpenR1-Math-220k(Face, 2025), Light-R1-SFT(Wen et al., 2025a), Bespoke-Stratos-17k(Labs, 2025), SYNTHETIC-1(Mattern et al., 2025), s1K-1.1(Muennighoff et al., 2025), LIMO(Ye et al., 2025) , leaf, fill=SFT3] ] [Supervised Fine-tuning Models (§2.2), level2,fill=SFT2 [e.g., OpenThinker(Team, 2025a), OpenR1-Qwen-7B(Face, 2025), Light-R1(Wen et al., 2025a), Bespoke-Stratos(Labs, 2025), SYNTHETIC-1-SFT(Mattern et al., 2025), s1.1(Muennighoff et al., 2025), LIMO(Ye et al., 2025) , leaf, fill=SFT3] ] ] [Reinforcement Learning
from Verifiable Rewards (§3), level1,fill=RL1 [Data Curation (§3.1), level2,fill=RL2 [e.g., Open-Reasoner-Zero(Hu et al., 2025), OWM(Gandhi et al., 2025), LIMR(Li et al., 2025b), DeepScaleR(Luo et al., 2025b), Skywork-OR1(He et al., 2025a), DAPO(Yu et al., 2025), Open-R1(Face, 2025), DeepCoder(Luo et al., 2025a) ,leaf,fill=RL3] ] [Algorithm Design (§3.2.1), level2,fill=RL2 [e.g., PPO(Schulman et al., 2017), GRPO(Shao et al., 2024), Dr. GRPO(Liu et al., 2025c), DAPO(Yu et al., 2025), VAPO(Yue et al., 2025b), VC-PPO(Yuan et al., 2025), CPPO(Lin et al., 2025b), REINFORCE(Ahmadian et al., 2024), REINFORCE++(Hu, 2025), REINFORCE-Kimi(Kimi-Team et al., 2025), GPG(Chu et al., 2025b) , leaf,fill=RL3] ] [Reward Curation (§3.2.2), level2,fill=RL2 [e.g., Logic-RL(Xie et al., 2025a), Oat-Zero(Liu et al., 2025b), L1(Aggarwal & Welleck, 2025), Efficient-Reasoning(Arora & Zanette, 2025), Open-Reasoner-Zero(Hu et al., 2025), Dr. GRPO(Liu et al., 2025c), DAPO(Yu et al., 2025), Demystify-Long-CoT(Yeo et al., 2025), Light-R1(Wen et al., 2025a) , leaf,fill=RL3] ] [Sampling Strategy (§3.2.3), level2,fill=RL2 [e.g., Light-R1(Wen et al., 2025a), Open-Reasoner-Zero(Hu et al., 2025), SRPO(Zhang et al., 2025b), DAPO(Yu et al., 2025), Skywork-OR1(He et al., 2025a) , leaf,fill=RL3] ] ] [Other Techniques (§4), level1,fill=Other1 [Process-Reward-based RL (§4.1), level2,fill=Other2 [e.g., IPR(Xiong et al., 2024), PRMBench(Song et al., 2025), FineReason(Chen et al., 2025) rStar-Math(Guan et al., 2025), PRIME(Cui et al., 2025) , leaf,fill=Other3] ] [Preference Optimization (§4.1), level2,fill=Other2 [e.g., EXAONE(Research et al., 2025), Light-R1(Wen et al., 2025a), DPO-VP(Tu et al., 2025), RedStar(Xu et al., 2025a), DPO-R1(Zhang et al., 2025a) , leaf,fill=Other3] ] ] ]
@@ 2 Supervised Fine-tuning for Reasoning Language Models
DeepSeek-R1 distilled models (Guo et al., 2025), e.g., the DeepSeek-R1-Distill-Qwen series, exhibit strong reasoning capabilities despite their smaller sizes. Since then, several works (Face, 2025; Team, 2025a; Wen et al., 2025a; Zhao et al., 2025a) have attempted to reproduce this reasoning ability in smaller models by applying Supervised Fine-Tuning (SFT) on their own curated datasets. These datasets typically consist of math or coding problems and, more importantly, include one or more validated Chain-of-thoughts (CoTs) from DeepSeek-R1. This section aims to provide a comprehensive overview of how these studies approach the reproduction of distilled reasoning models.
@@ 2.1.1 Data Collection and Curation Pipeline
After raw data collection, multiple rounds of filtering are typically employed to enhance quality. These include deduplication (e.g., via embedding similarity or n-gram), rejection sampling, and groun-truth verification. To ensure high coverage and data richness, many datasets explicitly emphasize difficulty and diversity during the selection process, often using heuristics or model pass rates to prioritize harder problems. For example, Light-R1 (Wen et al., 2025a) applies thresholding based on model correctness to form a challenging subset from a broader base. Further, most datasets rely on verified CoTs or solutions to ensure correctness and quality. Verification methods vary by domain. For instance, math problems are often validated by Math Verify (Kydlíček, 2024), coding questions through execution or unit tests, and general tasks by LLM judges. This combination of domain-aware validation and selective retention allows curators to distill high-quality reasoning traces that better support supervised fine-tuning.
@@ 2.1.2 Existing Dataset Details
OpenThoughts.
OpenThoughts [^1] https://huggingface.co/datasets/open-thoughts/OpenThoughts-114k (Team, 2025a) curates a synthetic reasoning dataset with 114k examples from several sources. It covers multiple domains, including math, science, coding, and puzzles. The CoTs are generated by DeepSeek-R1 and verified. In particular, they use an LLM as a judge to verify the ground-truth answers for math and puzzle problems, and rely on code execution and unit tests to validate coding problems.
Open-R1.
OpenR1-Math-220k [^2] https://huggingface.co/datasets/open-r1/OpenR1-Math-220k (Face, 2025) is a large-scale dataset for math reasoning tasks. Face (2025) collect 220k math problems from NuminaMath 1.5 (LI et al., 2024) and generate 2 to 4 CoTs for each problem using DeepSeek-R1. To ensure that each problem includes at least one correct answer, most of the CoTs are verified by Math Verify (Kydlíček, 2024), with Llama-3.3-70B-Instruct (Dubey et al., 2024) serving as a judge for 12% of the samples. Among the 220k problems, 94k are considered higher quality. According to Face (2025), the 94k subset achieves better performance in SFT, as the 131k extended problems may contain easier questions.
Light-R1.
Bespoke Stratos.
Labs (2025) curates a reasoning dataset [^4] https://huggingface.co/datasets/bespokelabs/Bespoke-Stratos-17k of 17k examples distilled from DeepSeek-R1, covering domains such as coding, math, science, and puzzles. They apply rejection sampling to eliminate reasoning traces with incorrect solutions. In particular, they use GPT-4o-mini (OpenAI, 2024) as a judge to filter out the traces with incorrect answers of math questions, increasing the proportion of retained examples from 25% to 73%, compared to a rule-based approach.
@@ 2.2 Training & Performance Comparison
Bespoke-Stratos (Labs, 2025) Qwen2.5-7 / 32B-Instruct 20.0 / 63.3 82.0 / 93.0
OpenThoughts (Team, 2025a) Qwen2.5-7 / 32B-Instruct 31.3 / 68.0 23.3 / 49.3 83.2 / 90.6
Open-R1 (Face, 2025) Qwen2.5-Math-7B-Instruct 36.7 40.0 90.6
@@
LIMO (Ye et al., 2025) Qwen2.5-32B-Instruct 57.1 44.5 94.8
Table 2: Summary of recent projects including initial checkpoints and their corresponding benchmark results. Resutls are taken from corresponding papers. Dashes (–) indicate unavailable results. Note that the reported Open-R1 performance is from the “default” split of the dataset.
@@
Although the technical report of DeepSeek-R1 does not mention the training hyperparameters of the distillation models, we aggregate this information from replication studies to better understand common training setups. For long-context tasks such as complex reasoning, the RoPE scaling factor ($\theta$) and maximum context length in the model configuration are often adjusted to enable extended context capabilities (Chen et al., 2023). For example, Open-R1 (Face, 2025) sets $\theta=300{,}000$ and the context length to 32,768 tokens. Commonly used learning rates include $1.0\times 10^{-5}$ and $5.0\times 10^{-5}$, with typical batch sizes of 96 or 128. Additionally, packing is usually employed to improve training efficiency (Wang et al., 2024a).
@@ 3.1 RL Datasets
Open-Reasoner-Zero (Hu et al., 2025) identifies three key aspects for data curation: quantity, diversity, and quality. Open-Reasoner-Zero collects a total of 129k training data, of which 72k are mainly cleaned from OpenR1-Math-220k (Face, 2025), and the rest are collected from various sources, including AIME (up to 2023), MATH, Numina-Math collection and Tulu3 MATH. It also leverages additional synthetic data using programmatic approaches to cover other reasoning domains. In controlling data quality, Open-Reasoner-Zero filters out samples with unverifiable formats, such as multiple-choice and proof-oriented problems. Also, it filters out non-English samples for better training stability and final model performance. Additionally, Open-Reasoner-Zero selects a challenging subset of 13k samples from the complete dataset using a intermediate model checkpoint during training, which is used to support the curriculum learning on difficulty of this work, aiming to address the shortcomings of the model and enhancing its performance on challenging scenarios.
@@
CuratedThoughts (Hochlehnert et al., 2025b) is curated from prevailing SFT datasets, filtering out improper samples to support stable RL training. It is collected from OpenR1-Math-220k (Face, 2025), OpenThoughts-114k and OpenThoughts-114k-Math (Team, 2025a), removing multi-part questions, questions asking for a proof, questions referring to figures or charts, and questions without a valid answer. DAPO-Math-17k (Yu et al., 2025) modifies the questions from the AoPS website (AoPS, 2025) so that the expected answer would always be an integer. The simple answers making them easy to parse to minimize errors from math verifiers, providing accurate reward signals during RL training. LIMR (Li et al., 2025b) proposes the learning impact measurement to filter a small amount of samples whose learning patterns complement the model’s overall performance trajectory, demonstrating that these samples tend to be more valuable for optimization.
@@ 3.2.1 Algorithms
As the most prevalent outcome-reward-based RL methods, PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) are widely used for fine-tuning LLMs. Interestingly, recent replication studies have introduced various modifications to these methods, tailoring them for specific purposes to enhance training effectiveness. We review several representative RL-based LLM fine-tuning algorithms, including REINFORCE (Ahmadian et al., 2024), PPO (Schulman et al., 2017), GRPO (Shao et al., 2024), and their variants (Liu et al., 2025c; Yu et al., 2025; Yuan et al., 2025; Hu, 2025; Kimi-Team et al., 2025; Lin et al., 2025b). Moreover, we outline the modification details in these methods along with their underlying motivations, aiming to provide a clear overview of the methodological advancements in outcome-reward-based RL training methods.

### v2

@@ 1 Introduction
Reasoning language models (RLMs) have emerged as a transformative advancement in the evolution of large language models (LLMs), such as OpenAI o-series (Jaech et al., 2024; OpenAI, 2025), DeepSeek-R1 (Guo et al., 2025), and QwQ series (Qwen-Team, 2024; 2025b). Unlike conventional LLMs that merely generate unstructured responses, these models incorporate an explicit chain-of-thought process, providing step-by-step reasoning that mimics human cognitive processes–such as invoking self-verification, reflection, and more. This shift quickly attracted attention of the LLM research community, as it meets the growing demand for better explainability in complex tasks like mathematical problem solving, code generation, and logical reasoning, as well as the pursuit of steadily increasing accuracy. The significance of RLMs lies in their potential to enhance the accuracy of language models’ response with trustful rationales. Rather than only providing answers, these models also reveal their thought process, which is invaluable for educational purposes, critical decision-making, and debugging AI reasoning errors. DeepSeek-R1 (Guo et al., 2025) is a prime example of this new generation. It leverages innovative training techniques such as supervised fine-tuning (SFT) and reinforcement learning from verifiable rewards (RLVR), allowing the model to develop powerful reasoning behaviors with an affordable amount of supervised data. Especially, by using methods like RLVR and incorporating cold-start data, it achieves performance comparable to prior models (e.g. OpenAI o1 (Jaech et al., 2024)) with relatively low training costs.
However, many critical details remain undisclosed which is required to replicate the reasoning performance and the behaviors of self-verification and reflection exhibited by DeepSeek-R1. Although the DeepSeek-R1 team has publicly released their solution as training the model with Group Relative Policy Optimization (GRPO, Shao et al. (2024)) and rule-based reward systems, the optimal design of the reinforcement learning algorithm and reward system remains underexplored. Moreover, the training data and configurations of the SFT and RLVR stages are not released, leaving the impact on model performance to be further examined. In response, many replication works have attempted to explore the optimal design for RLMs from various perspectives (see Figure 1), yet a comprehensive list and comparison of these works are still lacking.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, align=center, fill=theme!20, , [Training Reasoning Models, ver [Supervised Fine-tuning (§2), level1, fill=SFT1 [Supervised Fine-tuning Datasets (§2.1), level2,fill=SFT2 [e.g., OpenThoughts-114k(OpenThoughts-Team, 2025), OpenR1-Math-220k(HuggingFace, 2025), Light-R1-SFT(Wen et al., 2025a), Bespoke-Stratos-17k(Bespoke-Labs, 2025), SYNTHETIC-1(Mattern et al., 2025), s1K-1.1(Muennighoff et al., 2025), LIMO(Ye et al., 2025) , leaf, fill=SFT3] ] [Supervised Fine-tuning Models (§2.2), level2,fill=SFT2 [e.g., OpenThinker(OpenThoughts-Team, 2025), OpenR1-Qwen-7B(HuggingFace, 2025), Light-R1(Wen et al., 2025a), Bespoke-Stratos(Bespoke-Labs, 2025), SYNTHETIC-1-SFT(Mattern et al., 2025), s1.1(Muennighoff et al., 2025), LIMO(Ye et al., 2025) , leaf, fill=SFT3] ] ] [Reinforcement Learning
from Verifiable Rewards (§3), level1,fill=RL1 [Data Curation (§3.1), level2,fill=RL2 [e.g., Open-Reasoner-Zero(Hu et al., 2025), OWM(Gandhi et al., 2025), LIMR(Li et al., 2025b), DeepScaleR(Luo et al., 2025b), Skywork-OR1(He et al., 2025a), DAPO(Yu et al., 2025), Open-R1(HuggingFace, 2025), DeepCoder(Luo et al., 2025a) ,leaf,fill=RL3] ] [Algorithm Design (§3.2.1), level2,fill=RL2 [e.g., PPO(Schulman et al., 2017), GRPO(Shao et al., 2024), Dr. GRPO(Liu et al., 2025c), DAPO(Yu et al., 2025), VAPO(Yue et al., 2025b), VC-PPO(Yuan et al., 2025), CPPO(Lin et al., 2025b), REINFORCE(Ahmadian et al., 2024), REINFORCE++(Hu, 2025), REINFORCE-Kimi(Kimi-Team et al., 2025), GPG(Chu et al., 2025b) , leaf,fill=RL3] ] [Reward Curation (§3.2.2), level2,fill=RL2 [e.g., Logic-RL(Xie et al., 2025a), Oat-Zero(Liu et al., 2025b), L1(Aggarwal & Welleck, 2025), Efficient-Reasoning(Arora & Zanette, 2025), Open-Reasoner-Zero(Hu et al., 2025), Dr. GRPO(Liu et al., 2025c), DAPO(Yu et al., 2025), Demystify-Long-CoT(Yeo et al., 2025), Light-R1(Wen et al., 2025a) , leaf,fill=RL3] ] [Sampling Strategy (§3.2.3), level2,fill=RL2 [e.g., Light-R1(Wen et al., 2025a), Open-Reasoner-Zero(Hu et al., 2025), SRPO(Zhang et al., 2025b), DAPO(Yu et al., 2025), Skywork-OR1(He et al., 2025a) , leaf,fill=RL3] ] ] [Other Techniques (§4), level1,fill=Other1 [Process-Reward-based RL (§4.1), level2,fill=Other2 [e.g., IPR(Xiong et al., 2024), PRMBench(Song et al., 2025), FineReason(Chen et al., 2025) rStar-Math(Guan et al., 2025), PRIME(Cui et al., 2025) , leaf,fill=Other3] ] [Preference Optimization (§4.1), level2,fill=Other2 [e.g., EXAONE(LG-Research et al., 2025), Light-R1(Wen et al., 2025a), DPO-VP(Tu et al., 2025), RedStar(Xu et al., 2025a), DPO-R1(Zhang et al., 2025a) , leaf,fill=Other3] ] ] ]
@@ 2 Supervised Fine-tuning for Reasoning Language Models
DeepSeek-R1 distilled models (Guo et al., 2025), e.g., the DeepSeek-R1-Distill-Qwen series, exhibit strong reasoning capabilities despite their smaller sizes. Since then, several works (HuggingFace, 2025; OpenThoughts-Team, 2025; Wen et al., 2025a; Zhao et al., 2025a) have attempted to reproduce this reasoning ability in smaller models by applying Supervised Fine-Tuning (SFT) on their own curated datasets. These datasets typically consist of math or coding problems and, more importantly, include one or more validated Chain-of-thoughts (CoTs) from DeepSeek-R1. This section aims to provide a comprehensive overview of how these studies approach the reproduction of distilled reasoning models.
@@ 2.1.1 Data Collection and Curation Pipeline
After raw data collection, multiple rounds of filtering are typically employed to enhance quality. These include deduplication (e.g., via embedding similarity or n-gram), rejection sampling, and ground-truth verification. To ensure high coverage and data richness, many datasets explicitly emphasize difficulty and diversity during the selection process, often using heuristics or model pass rates to prioritize harder problems. For example, Light-R1 (Wen et al., 2025a) applies thresholding based on model correctness to form a challenging subset from a broader base. Further, most datasets rely on verified CoTs or solutions to ensure correctness and quality. Verification methods vary by domain. For instance, math problems are often validated by Math Verify (Kydlíček, 2024), coding questions through execution or unit tests, and general tasks by LLM judges. This combination of domain-aware validation and selective retention allows curators to distill high-quality reasoning traces that better support supervised fine-tuning.
@@ 2.1.2 Existing Dataset Details
OpenThoughts.
OpenThoughts [^3] https://huggingface.co/datasets/open-thoughts/OpenThoughts-114k (OpenThoughts-Team, 2025) curates a synthetic reasoning dataset with 114k examples from several sources. It covers multiple domains, including math, science, coding, and puzzles. The CoTs are generated by DeepSeek-R1 and verified. In particular, they use an LLM as a judge to verify the ground-truth answers for math and puzzle problems, and rely on code execution and unit tests to validate coding problems.
Bespoke Stratos.
Bespoke-Labs (2025) curates a reasoning dataset [^4] https://huggingface.co/datasets/bespokelabs/Bespoke-Stratos-17k of 17k examples distilled from DeepSeek-R1, covering domains such as coding, math, science, and puzzles. They apply rejection sampling to eliminate reasoning traces with incorrect solutions. In particular, they use GPT-4o-mini (OpenAI, 2024) as a judge to filter out the traces with incorrect answers of math questions, increasing the proportion of retained examples from 25% to 73%, compared to a rule-based approach.
Light-R1.
Open-R1.
OpenR1-Math-220k [^6] https://huggingface.co/datasets/open-r1/OpenR1-Math-220k (HuggingFace, 2025) is a large-scale dataset for math reasoning tasks. HuggingFace (2025) collects 220k math problems from NuminaMath 1.5 (LI et al., 2024) and generate 2 to 4 CoTs for each problem using DeepSeek-R1. To ensure that each problem includes at least one correct answer, most of the CoTs are verified by Math Verify (Kydlíček, 2024), with Llama-3.3-70B-Instruct (Dubey et al., 2024) serving as a judge for 12% of the samples. Among the 220k problems, 94k are considered higher quality. According to HuggingFace (2025), the 94k subset achieves better performance in SFT, as the 131k extended problems may contain easier questions.
@@ 2.2 Training & Performance Comparison
OpenThoughts (OpenThoughts-Team, 2025) Qwen2.5-7 / 32B-Instruct 31.3 / 68.0 23.3 / 49.3 83.2 / 90.6
Bespoke-Stratos (Bespoke-Labs, 2025) Qwen2.5-7 / 32B-Instruct 20.0 / 63.3 82.0 / 93.0
Open-R1 (HuggingFace, 2025) Qwen2.5-Math-7B-Instruct 36.7 40.0 90.6
@@
Table 2: Summary of recent projects including initial checkpoints and their corresponding benchmark results. Results are taken from corresponding papers. Dashes (–) indicate unavailable results. Note that the reported Open-R1 performance is from the “default” split of the dataset.
@@
Although the technical report of DeepSeek-R1 does not mention the training hyperparameters of the distillation models, we aggregate this information from replication studies to better understand common training setups. For long-context tasks such as complex reasoning, the RoPE scaling factor ($\theta$) and maximum context length in the model configuration are often adjusted to enable extended context capabilities (Chen et al., 2023). For example, Open-R1 (HuggingFace, 2025) sets $\theta=300{,}000$ and the context length to 32,768 tokens. Commonly used learning rates include $1.0\times 10^{-5}$ and $5.0\times 10^{-5}$, with typical batch sizes of 96 or 128. Additionally, packing is usually employed to improve training efficiency (Wang et al., 2024a).
@@ 3.1 RL Datasets
Open-Reasoner-Zero (Hu et al., 2025) identifies three key aspects for data curation: quantity, diversity, and quality. Open-Reasoner-Zero collects a total of 129k training data, of which 72k are mainly cleaned from OpenR1-Math-220k (HuggingFace, 2025), and the rest are collected from various sources, including AIME (up to 2023), MATH, Numina-Math collection and Tulu3 MATH. It also leverages additional synthetic data using programmatic approaches to cover other reasoning domains. In controlling data quality, Open-Reasoner-Zero filters out samples with unverifiable formats, such as multiple-choice and proof-oriented problems. Also, it filters out non-English samples for better training stability and final model performance. Additionally, Open-Reasoner-Zero selects a challenging subset of 13k samples from the complete dataset using a intermediate model checkpoint during training, which is used to support the curriculum learning on difficulty of this work, aiming to address the shortcomings of the model and enhancing its performance on challenging scenarios.
@@
CuratedThoughts (Hochlehnert et al., 2025b) is curated from prevailing SFT datasets, filtering out improper samples to support stable RL training. It is collected from OpenR1-Math-220k (HuggingFace, 2025), OpenThoughts-114k and OpenThoughts-114k-Math (OpenThoughts-Team, 2025), removing multi-part questions, questions asking for a proof, questions referring to figures or charts, and questions without a valid answer. DAPO-Math-17k (Yu et al., 2025) modifies the questions from the AoPS website (AoPS, 2025) so that the expected answer would always be an integer. The simple answers making them easy to parse to minimize errors from math verifiers, providing accurate reward signals during RL training. LIMR (Li et al., 2025b) proposes the learning impact measurement to filter a small amount of samples whose learning patterns complement the model’s overall performance trajectory, demonstrating that these samples tend to be more valuable for optimization.
@@ 3.2.1 Algorithms
As the most prevalent outcome-reward-based RL methods, PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) are widely used for fine-tuning LLMs. Interestingly, recent replication studies have introduced various modifications to these methods, tailoring them for specific purposes to enhance training effectiveness. We review several representative RL-based LLM fine-tuning algorithms, including REINFORCE (Ahmadian et al., 2024), PPO (Schulman et al., 2017), GRPO (Shao et al., 2024), and their variants (Liu et al., 2025c; Yu et al., 2025; Yuan et al., 2025; Hu, 2025; Kimi-Team et al., 2025; Lin et al., 2025b). Moreover, we outline the modifications in these methods along with their underlying motivations, aiming to provide a clear overview of the methodological advancements in outcome-reward-based RL training methods.